The cheetah can maintain it's top speed for only about 20 s

The fastest land animal is the cheetah, which runs at a speed of up to 101 km/h. The second-fastest is the antelope, which runs at a speed of up to 88 km/h. (a). Suppose that a cheetah begins to chase an antelope. If the antelope has a head start of 40 m, how long does it take the cheetah to catch the antelope? How far will the cheetah have traveled by this time?
Answer #1

Vrel =101-88=13 km /hr

Time to catch if run at const speed

T = 0.04*3600/13 = 11.08 sec

Distance travel = VxT = 101*11.08*1000/3600 = 310.7 mtr